OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 8 π A = 4√2 16 64

OpenStudy (hysusonic):

whats \[\pi r^2\]

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ C=2\pi r }\) \(\LARGE\color{black}{ 8\pi=2\pi r }\) \(\LARGE\color{black}{ 8=2 r }\) \(\LARGE\color{black}{ 4= r }\) then use, \(\LARGE\color{blue}{ A=\pi r^2 }\)

OpenStudy (anonymous):

thnx:)

OpenStudy (solomonzelman):

anytime:)